OpenStudy (anonymous):

this question say write each question in standard form and please show work y-5=3/2(x+4)

OpenStudy (anonymous):

is x+4 in denominator?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

is it y-5=3(x+4)/2

OpenStudy (anonymous):

you didnt show the steps how did you get it

OpenStudy (anonymous):

i didn't solve it yet, is this the correct form you say? y-5=3(x+4)/2

OpenStudy (anonymous):

Standard form for a line is the following:\[ax+by=c\]You need to do whatever's necessary to your equation to get it into that form. You need to get rid of fractions, so multiply both sides by 2 first:\[2(y-5)=3(x+4)\]and expand\[2y-10=3x+12\]Now subtract 2y from both sides\[-10=3x-2y+12\]and subtract 12 from both sides\[-22=3x-2y\]So your equation in standard form is\[3x-2y=-22\]